Accounting question, thank you for the help!At January 1, 2021, Café Med leased restaurant equipment from Crescent Corporation under a nine-year lease agreement. The lea
Answer #1

Solution-

1. Effect on earnings $(29000)
2. Lease payable balance (end of year) $132349
Right-of-use asset balance (end of year) $161349

Working-1

Present value of annuity = 5.96764 (n=9 , i=12%)

29000 * 5.96764 =173061.56(right of use)

12%(173062-29000) =17287.44(interest exp)

29000 - 17287 = 11713(amortization-right of use)

11713+17287 = 29000

Working-2

173062 - 29000 - 11713 = 132349 (end of yr bal-liab)

Working-3

173062 - 11713 = 161349 (end of yr bal-assets)